OpenStudy (anonymous):

Any help would be great. Or just the formulas. Total Surface area of a Square Prism with each side of the base equal to 10 in. and a height of 9 in. Needs to be expressed in cubic inches. Thank you.

OpenStudy (automaticloveletter):

v=a^2*h

OpenStudy (anonymous):

I'm looking for the Total surface area.

OpenStudy (automaticloveletter):

Oh sorry xDD

OpenStudy (automaticloveletter):

A=2*a^2+4*a*h

OpenStudy (anonymous):

2*10^2+4*10*9=560 cubic inches?